OpenStudy (anonymous):

Help please! Part A: Solve A = 9 over 2(x + 28) for x. (4 points) Part B: Determine the value of x when A = 135. (2 points) Part C: Solve –np – 40 > 10 for n. Show your work. (4 points)

OpenStudy (here_to_help15):

help is here : P

OpenStudy (here_to_help15):

9/2(x+28) = A multiply by the recprocal 2A/9 = x + 28 subtract 28 2A/9 - 28 = x

OpenStudy (here_to_help15):

B. Just evaluate it, A=135 put that in and solve for x. C. –np – 40 > 10 add 40 -np > 50 solving for n, divide by -p and flip the sign n < 50/-p

OpenStudy (here_to_help15):

you still with me?

OpenStudy (here_to_help15):

??

OpenStudy (anonymous):

Yes, sorry! Thank you so much. It helped alot

OpenStudy (antonio_xx2):

@Squirrels

OpenStudy (squirrels):

So I assume you just need help with C since he already helped you with A and B right?